OpenStudy (anonymous):

Four men working together can dig a ditch in 42 days. They begin the work together but one man works only for half of the days. How much time will be taken to complete the job?

OpenStudy (whpalmer4):

Figure out the fraction of the whole job one man working for one day will complete.

OpenStudy (anonymous):

i men =1/168 1/2 men=1/336,

OpenStudy (whpalmer4):

Then, multiply that by 3.5 to get the amount of e job done in one average day by the 4 men.

OpenStudy (anonymous):

48 days

OpenStudy (whpalmer4):

That's what I get, too.

OpenStudy (anonymous):

4*42/3.5=48

OpenStudy (dumbcow):

assuming each man works at same rate R 4R = 1/42 --> R = 1/168 it takes 1 man 168 days to dig the ditch Let the num of days be N for N/2 days, all four men work together at rate of 1/42 percent of ditch = rate*time --> N/2 * 1/42 = N/84 for the next N/2 days only 3 men work at a rate of 3/168 --> N/2 * 3/168 = 3N/336 = N/112 all together 1 ditch is dug add and set equal to 1 --> N/84 + N/112 = 1 \[N = \frac{1}{\frac{1}{84}+\frac{1}{112}}\] using calculator :) N = 48

OpenStudy (whpalmer4):

Figuring out the rate is the key to solving these equations.

OpenStudy (dumbcow):

oh crap i did it the long way :{

OpenStudy (whpalmer4):

But you provided a valuable check on our result, so it is not unappreciated :-)

OpenStudy (anonymous):

dumbcow came first so i am giving the medal to him.sry whpalmer4 not giving the medal .becoz u said the easiest way

OpenStudy (whpalmer4):

How exactly did dumbcow come first? Just curious, and don't really care about the medal...

OpenStudy (anonymous):

he started answering the question before u came so i said @whpalmer4 .dont mistake me

OpenStudy (anonymous):

understood?